i need help in solving rational exponets like this (1/16) TO THE 1/4 POWER how do you solve that
Answer by lwsshak3(11628)

i need help in solving rational exponets like this (1/16) TO THE 1/4 POWER how do you solve that
**
The simpliest way to do this is to raise both numerator and denominator to the 1/4 power
Raising the number 1 to any power=1, so 1^1/4=1
Raising 16 to the 1/4 power is like taking the fourth root of 16,. that is, 16^(1/4)=2
So (1/16)^(1/4)=1/2
..
Another way to show it:
(1/16)^(1/4)= (1/2^4)^(1/4)=(2^-4)^(1/4)=2^-1=1/2